# Fast Polynomial Root-Finding and Approximation of the Eigenvalues of a Matrix

### v36

@@ 13 Background
13.1 Definitions
We recall or redefine some concepts from the Introduction and Part II.
  • ${\rm NR}_{t}(x):=\frac{t(x)}{t^{\prime}(x)}$ is Newton’s ratio, ${\rm NIR}_{t}(x):=\frac{t^{\prime}(x)}{t(x)}$ is Newton’s inverse ratio for a polynomial $t(x)$. We write NR$(x):={\rm NR}_{p}(x)$ and NIR$(x):={\rm NIR}_{p}(x)$ (cf. (3.2)).
@@ 13.1 Definitions
  • $\Delta(\mathbb{S})$, $CH(\mathbb{S})$, $X(\mathbb{S})$ and #$\#(\mathbb{S})$ denote the diameter, root set, convex hull, and index (root set’s cardinality) of a set $\mathbb{S}$ on the complex plane, respectively.
@@
    We say isolated for “$\theta$-isolated" if $\theta-1$ exceeds a positive constant.
@@
    We say rigid for “$\eta$-rigid" if $\eta$ exceeds a positive constant.
  • $|u|:=||{\bf u}||_{1}:=\sum_{i=0}^{q-1}|u_{i}|$, $||{\bf u}||:=||{\bf u}||_{2}:=(\sum_{i=0}^{q-1}|u_{i}|^{2})^{\frac{1}{2}}$, and $|u|_{\rm max}:=||{\bf u}||_{\infty}:=\max_{i=0}^{q-1}|u_{i}|$ denote the 1-, 2-, and $\infty$-norm, respectively, of a polynomial $u(x)=\sum_{i=0}^{q-1}u_{i}x^{i}$ and its coefficient vector [^22] Here and hereafter ${\bf v}$ denotes a column vector, while ${\bf v}^{T}$ denotes its transpose. ${\bf u}=(u_{i})_{i=0}^{d-1}$, and then (see [48, Eqs. (2.2.5)–(2.2.7)])
@@
  • Generalize Def. 2.1 of $\sigma$-soft e/i tests as follows: For $p(x)$ of (1.1), $\sigma>1$, and integers $\ell$ and $m$ such that $1\leq\ell\leq m\leq d$ and #$\#(D(0,1))\leq m$, a $\sigma$-soft $\ell$-test, or just $\ell$-test for short (in particular e/i test turns into 1-test), either outputs 1 and stops if it detects that $r_{d-\ell+1}\leq\sigma$, that is, #$\#(D(0,\sigma))\geq\ell$, or outputs 0 and stops if it detects that $r_{d-\ell+1}>1$, that is, [^23] Bounds $r_{d-\ell+1}\leq\sigma$ and $r_{d-\ell+1}>1$ can hold simultaneously (cf. Fig. 1b), but as soon as an $\ell$-test verifies any of them, it stops without checking if the other bound also holds. #$\#(D(0,1))<\ell$. An $\ell$-test applied to the polynomial $t(y)$ of (1.3) is said to be an $\ell$-test for the disc $D(c,\rho)$ and/or the circle $C(c,\rho)$ and also to be an $\ell$-testc,ρ. [^24] We will only use $\ell$-tests for $\ell=1$ (which are e/i tests) and $\ell=m$.
Observation 13.1.
@@ 13.2 Reverse polynomial
Immediately observe the following properties of the reverse polynomial of (1.5).
@@
13.3 Root-squaring and root-lifting
The DLG root-squaring iterations (3.7) by Dandelin 1826, Lobachevsky 1834, and Gräffe 1837 compute the polynomials $p_{2^{i}}(x):=\prod_{j=1}^{d}(x-z_{j}^{2^{i}})$, for $i=0,1,\dots$. By applying $h$-lifting of (3.6) we extend the definition of $p_{h}(x)$ to all $h\in\mathbb{N}$.
As we noticed in the introduction, the known studies of DLG iterations always focus and stumble on numerically unstable computation of the coefficients of the polynomials $p_{i}(x)$ [86, 113], but we only compute their values bypassing those problems.
@@ 13.3 Root-squaring and root-lifting
${\rm NIR}_{p_{h}}(x)=\frac{1}{{\rm NR}_{p_{h}}(x)}\approx\frac{p_{h}(x)-p_{h}(x-\delta)}{p_{h}(x)\delta}=\frac{p_{h}^{\prime}(y)}{p_{h}(x)}.$ (13.3)
@@
13.4 The distance to a closest root and fast inclusion tests
Next we estimate the distance from the origin to the closest root. The distance to the farthest root $z_{1}$ is given by the same estimates applied to the reverse polynomial $p_{\rm rev}(x)$.
For bounds $r_{d}$ and $r_{1}$ in terms of the coefficients of $p$, see [80, Sec. 1.10]; in particular according to Fujihara’a bounds of [41]:
@@ 13.4 The distance to a closest root and fast inclusion tests
For a complex $c$ and positive integers $j$ and $h$, Eqn. (3.1) and Observation 13.2 combined imply the following relationships independent of $p_{d}$:
$r_{d}(c)\leq d~{}|{\rm NR}(c)|,$ (13.4)
$r^{h}_{j}(0,p)=r_{j}(0,p_{h}).$ (13.5)
@@
We can assume that $|z_{j}|>\theta$ if and only if $j>m$ and then deduce from Eqn. (3.1) that $|{\rm NIR}(c)|\leq\sum_{j=1}^{m}\frac{1}{|c-z_{j}|}+\sum_{j=m+1}^{d}\frac{1}{|c-z_{j}|}<\frac{m}{r_{d}(c)}+\frac{d-m}{\theta-1}$. ∎
Certify inclusion into the disc $D(0,\sigma)$ if $\rho_{d}(0,p_{h})\leq\sigma^{h}$, $\sigma>1$ and $h>0$, although this bound is very poor for worst case polynomials $p$:
@@
Next we argue that such deficiency only occurs for a narrow class of polynomials $p$ and comment on some heuristic recipes to counter it.
Notice that $\frac{1}{r_{d}(c)}\leq|{\rm NIR}(c)|=\frac{1}{d}\Big{|}\sum_{j=1}^{d}\frac{1}{c-z_{j}}\Big{|}$ by virtue of (3.1), and so approximation to the root radius $r_{d}(c,p)$ is poor if and only if severe cancellation occurs in the summation of the $d$ roots.
Formal probabilistic support for this observation is hard, but intuitively such a cancellation only occurs for a very narrow class of polynomials $p$. If so, the above non-costly test certifies inclusion for a very large class of polynomials $p$ [^26] One can try some empirical recipes. Clearly, the problem persists for the root radius $r_{d}(w)$ where $p^{\prime}(w)$ vanishes; rotation of the variable $p(x)\leftarrow t(x)=p(ax)$ for $|a|=1$ does not fix it, but the bound $r_{d}\leq|p_{0}/p_{d}|^{1/d}$ and shifts $p(x)\leftarrow t(x)=p(x-c)$ for $c\neq 0$ combined should fix it. We must also study the case where e/i tests output exclusion, however, because this occurs in more than 25% of all e/i tests in a subdivision process:
@@
Notice that $\sigma_{i+1}=4(\sigma_{i}-d_{i}-f_{i})~{}{\rm for}~{}i=1,2,\dots,t+1$, sum these equations for all $i$, substitute $\sigma_{1}=4$ and $\sigma_{t+1}=0$, and obtain $3\Sigma+4\geq 4D+4F$. Represent subdivision process by a tree with $F\leq m$ leaves whose other nodes are components made up of suspect squares, to obtain pessimistic lower bounds $\Sigma+1\geq 2m\geq 2F$ and hence $D\geq\Sigma/4+1/2$. ∎
@@ 14.1 Overview
Map (1.3) and Cor. 3.1 together reduce an e/i test for any disc to the test for the unit disc $D=D(0,1)$, and we devise soft e/i tests for that disc assuming that #$\#(D(0,2))\leq m$ for a fixed $m$, $1\leq m\leq d$. Our Alg. 14.1 of Sec. 14.2 and Alg. 14.2 of Sec. 14.3 are deterministic specifications of Alg. 3.1, which is always correct when it claims inclusion, by virtue of Lemma 3.1. We prove that both of these specifications are also correct when they output exclusion. Namely, Algs. 14.1 and 14.2 evaluate ${\rm NIR}(x)$ at $hq$ equally spaced points of the unit circle $C(0,1)$ for $h=O(\log(d))$ and correctly claim exclusion if at all these points $|{\rm NIR}(x)|\leq\frac{1}{2\sqrt{q}}$ for $q>d$ or if $|{\rm NIR}(x)|\leq\frac{1}{3\sqrt{q}}$ for $q>m$, respectively. Our Example 14.1 shows that we cannot devise such deterministic algorithms for $q$ decreased below $m$, but in Sec. 14.4 we devise Las Vegas randomized e/i tests that evaluate polynomial $p(x)$ at $O(\log(d))$ points $x$ overall under the Random Root Model.
@@ 14.2 Basic deterministic e/i test
COMPUTATIONS: For $g=0,1,\dots,q-1$ recursively compute
@@
Alg. 14.1 outputs $\sigma$-soft inclusion correctly by virtue of Lemma 3.1. Next we prove that it also correctly claims exclusion.
@@
Now, if Alg. 14.1 outputs exclusion, then $1+\frac{d}{v_{g}}>\sigma^{h}$ for $g=0,1,\dots,q-1$, where $\sigma^{h}>1+2d\sqrt{q}$ according to initialization of Alg. 14.1. Hence $\frac{1}{v_{g}}>2\sqrt{q}$, and so the assumptions of Prop. 14.1 hold for $p_{h}(x)$ replacing $p(x)$. Therefore, the theorem implies that the disc $D(0,1)$ contains no zeros of $p_{h}(x)$, but then it contains no zeros of $p(x)$ as well (see Observation 13.2 (iii)), and so exclusion is correct.
@@
Alg. 14.1 is a $\sigma$-soft e/i test, certifying inclusion at the cost of evaluation of polynomial $p(x)$ at $2hj$ points $x$ or exclusion at the cost of evaluation at $2hq$ points $x$ for three integers $h$, $j$, and $q$ such that $1\leq j\leq q$, $q>m$, and $h:=\lceil\log_{\sigma}(1+2d\sqrt{q})\rceil$, $h=O(\log(d))$ for $\sigma-1$ exceeding a positive constant.
@@ 14.3 A faster deterministic e/i test
Next we supply some relevant estimates.
@@
Prop. 13.2, for $p_{h}$ replacing $p$, and the assumed bound $\frac{d-m}{\theta^{h}-1}\leq\frac{1}{6\sqrt{q}}$ together imply
@@
Alg. 14.2 claims inclusion if $v_{g}\geq\frac{1}{3\sqrt{q}}$ for some $g$, but then
@@
Now Observation 13.2 (iii) implies that Alg. 14.2 claims inclusion correctly.
Next prove that it also claims exclusion correctly.
@@
Combine these bounds to obtain $\max_{g=0}^{q-1}\Big{|}\frac{f^{\prime}(\zeta_{q}^{g})}{f(\zeta_{q}^{g})}\Big{|}<\frac{1}{2\sqrt{q}}$. To complete the proof, apply Prop. 14.1 for $f(x)$ replacing $p(x)$ and for $m$ replacing $d$. ∎
Readily extend Prop. 14.2 to Alg. 14.2 by decreasing its bounds on $q$ and $h$ provided that #$\#(D(0,\theta))=m$ for $\theta\geq 1+6(d-m)\sqrt{q}$.
@@
Alg. 14.2 is a $\sigma$-soft e/i test: it certifies either inclusion or exclusion at the cost of evaluation of $p(x)$ at $2hj$ or $2hq$ points $x$, respectively, for three integers $h$, $j$, and $q$ such that $1\leq j\leq q$, $q>m:=\#(D(0,\theta)),~{}\sigma^{h}>1+6m\sqrt{q}$, $\theta\geq 1+6(d-m)\sqrt{q}$, and $h:=\lceil\log_{\sigma}(1+6m\sqrt{q})\rceil$, so that $h=O(\log(m))$ if $\sigma-1$ exceeds a positive constant.
@@ 14.4 A randomized e/i test
14.4.1. Introductory comments. Would Algs. 14.1 and 14.2 stay correct if we apply them at fewer evaluation points? They have consistently stayed correct for $q$ of order $\log^{2}(d)$ in extensive numerical tests of [60, 63] applied to $p(x)$ rather than to $p_{i}(x)$ for $i>0$. The tests reported exclusion where the values $|s_{h,q}|$ were reasonably small for $s_{h,q}:=\frac{1}{q}\sum_{g=0}^{q-1}\zeta^{(h+1)g}~{}\frac{p^{\prime}(\zeta^{g})}{p(\zeta^{g})},~{}{\rm for}~{}h=0,1,2$ (cf. Eqn. 7.2 in Part II), and reported inclusion otherwise. Clearly, $|s_{h,q}|\leq\delta$ for all $h$ if $|p(\zeta^{g})|\leq\delta/q$ for $g=0,1,\dots,q-1$, while inclusion is certified unless $|p(\zeta^{g})|$ is small for all $g$.
@@
Clearly, [60, 63] have never encountered worst case inputs in their extensive tests. Next we formally support this observation by introducing probability structure in the space of polynomials $p(x)$ and $t(x)=p(\frac{x-c}{\rho})$. Shifts destroy all customary probability distributions in the space of the coefficients of $p(x)$, but we prove that our e/i tests that evaluate $p(x)$ at $O(\log(d)$ points are correct whp under the Random Root Model.
@@
Hence $|z-\bar{z}|=O(1/\beta^{g})$ for a fixed $\beta>1$. Combine this bound with (14.6) and (14.7), choosing sufficiently large $g$ of order $\log d$ unless $\sigma|\bar{z}|<\beta$.
@@
For a fixed $\sigma>1$, e.g., for $\sigma=1.2$, choose a constant $\beta>1$ and sufficiently large $g=O(\log(d))$ such that $\gamma<\pi$ in (14.4). Then under the Random Root Model the probability $P$ of the output error for Alg. 3.1(rrm) is less than a constant $\nabla<1$. We can decrease the error probability below $\nabla^{v}$ by applying the algorithm $v$ times, in particular below $d^{a}$ for any fixed positive constant $a$ and for $v=O(\log(d))$.
Based on Cor. 2.2 we reduce classical subdivision root-finder to performing $O(mb)$ soft e/i tests, but in Part IV we decrease that bound to $w=O(m\log(b))$ tests. These tests magnify the error bounds by at most a factor of $w$. Since $m\leq d$, we can ensure correctness of the resulting e/i test whp provided that $\log(b)=o(d^{a-1})$. We arrive at the following
@@
Under Random Root Model one only need to apply a subdivision root-finder that involves $O(m\log(b)))$ soft Las Vegas randomized e/i tests to ensure the output error bound $\epsilon=1/2^{b}$ whp provided that $\log(b)=d^{O(1)}$ and that every test is reduced to $v=O(\log(d))$ applications of Alg. 3.1(rrm).
@@
It is tempting to combine root-lifting or root-squaring with algorithms of [115], which approximate Cauchy integral (3.4), to obtain alternative randomized e/i test supporting Cor. 14.1, but for that we would have to raise disc isolation from order of $1+1/m$ to a constant, and this would require too costly lifting of order $m$.

### v37

@@ 6 Background
6.1 Basic definitions
We extend the list of definitions used in the Introduction.
  • ${\rm NR}_{t}(x):=\frac{t(x)}{t^{\prime}(x)}$ and ${\rm NIR}_{t}(x):=\frac{t^{\prime}(x)}{t(x)}$ are Newton’s ratio and Newton’s inverse ratio, respectively, for a polynomial $t(x)$. We write NR$(x):={\rm NR}_{p}(x)$ and NIR$(x):={\rm NIR}_{p}(x)$ (cf. (3.1)).
@@ 6.1 Basic definitions
  • $\Delta(\mathbb{S})$, $MCD(\mathbb{S})$, $X(\mathbb{S})$, and #$\#(\mathbb{S})$ denote the diameter, root set, minimal covering disc, and index (root set’s cardinality) of a set $\mathbb{S}$ on the complex plane, respectively.
@@
    We say “isolated" for “$\theta$-isolated" if $\theta-1$ exceeds a positive constant.
@@
    We say “rigid" for “$\eta$-rigid" if $\eta$ exceeds a positive constant.
  • $|u|:=||{\bf u}||_{1}:=\sum_{i=0}^{q-1}|u_{i}|$, $||{\bf u}||:=||{\bf u}||_{2}:=(\sum_{i=0}^{q-1}|u_{i}|^{2})^{\frac{1}{2}}$, and $|u|_{\rm max}:=||{\bf u}||_{\infty}:=\max_{i=0}^{q-1}|u_{i}|$ denote the 1-, 2-, and $\infty$-norm, respectively, of a polynomial $u(x)=\sum_{i=0}^{q-1}u_{i}x^{i}$ and its coefficient vector ${\bf u}=(u_{i})_{i=0}^{d-1}$, and then (see [35, Eqs. (2.2.5)–(2.2.7)])
@@
Observation 6.1.
@@ 6.2 Reverse polynomial
Immediately observe the following properties of the reverse polynomial (1.5).
@@
6.3 Root-lifting with divided differences
@@
${\rm NIR}_{p_{h}}(x)\approx\frac{p_{h}(x)-p_{h}(x-\delta)}{p_{h}(x)\delta}=\frac{p_{h}^{\prime}(y)}{p_{h}(x)}.$ (6.3)
@@
6.4 The distance to a closest root and certified inclusion
[63, Sec. 1.10] lists various bounds on $r_{d}$ and $r_{1}$ in terms of the coefficients of $p$, in particular [31]:
@@ 6.4 The distance to a closest root and certified inclusion
For a complex $c$ and positive integers $j$ and $h$, Eqn. (3.1) and Observation 6.2 combined imply that
$r_{d}(c)\leq d~{}|{\rm NR}(c)|,~{}{\rm and}~{}r^{h}_{j}(0,p)=r_{j}(0,p_{h}).$ (6.4)
@@
Assume that $|z_{j}|>\theta$ if and only if $j>m$ and then deduce from Eqn. (3.1) that $|{\rm NIR}(c)|\leq\sum_{j=1}^{m}\frac{1}{|c-z_{j}|}+\sum_{j=m+1}^{d}\frac{1}{|c-z_{j}|}<\frac{m}{r_{d}(c)}+\frac{d-m}{\theta-1}$. ∎
Certify inclusion into the disc $D(0,\sigma)$ if $\rho_{d}(0,p_{h})\leq\sigma^{h}$, $\sigma>1$ and $h>0$, although this bound is extremely poor for worst case polynomials $p$:
@@
Notice that $\frac{1}{r_{d}(c)}\leq|{\rm NIR}(c)|=\frac{1}{d}\Big{|}\sum_{j=1}^{d}\frac{1}{c-z_{j}}\Big{|}$ by virtue of (3.1), and so approximation to the root radius $r_{d}(c,p)$ is poor if and only if severe cancellation occurs in the summation of the $d$ roots. It seems that such a cancellation only occurs for a very narrow class of polynomials $p$, and if so, then the above non-costly test certifies inclusion for a very large class of polynomials $p$ [^15] Shifts of the variable $p(x)\leftarrow t(x)=p(x-c)$ for random $c$ should further enlarge this class.
6.5 Subdivision root-finding involves exclusion
It is not enough for us to have non-costly certification of inclusion because exclusion occurs in more than 25% of all e/i tests in a subdivision process:
@@ 6.5 Subdivision root-finding involves exclusion
Notice that $\sigma_{i+1}=4(\sigma_{i}-d_{i}-f_{i})~{}{\rm for}~{}i=1,2,\dots,t+1$, sum these equations for all $i$, substitute $\sigma_{1}=4$ and $\sigma_{t+1}=0$, and obtain $3\Sigma+4\geq 4D+4F$. Represent subdivision process by a tree with at most $F\leq m$ leaves whose other nodes are components made up of suspect squares, to obtain pessimistic lower bounds $\Sigma+1\geq 2m\geq 2F$ and hence $D\geq\Sigma/4+1/2$. ∎
@@ 7.1 Overview
Our Alg. 7.1 of Sec. 7.2 and Alg. 7.2 of Sec. 7.3 are deterministic specifications of Alg. 3.1. By virtue of Lemma 3.1 all of them are correct when they claim inclusion, but we will prove that they are also correct when output exclusion. Namely, Algs. 7.1 and 7.2 evaluate ${\rm NIR}(x)$ at $hq$ equally spaced points of the unit circle $C(0,1)$ for $h=O(\log(d))$ and certify exclusion if at all these points $|{\rm NIR}_{p_{h}}(x)|\leq\frac{1}{2\sqrt{q}}$ for $q>d$ or if $|{\rm NIR}_{p_{h}}(x)|\leq\frac{1}{3\sqrt{q}}$ for $q>m$, respectively, and otherwise certify inclusion. Our Example 7.1 shows that we cannot devise such deterministic algorithms for $q<m$, but our Las Vegas e/i test evaluates polynomial $p(x)$ at $O(\log(d))$ points $x$ overall under the Random Root Model.
@@ 7.2 Basic deterministic e/i test
COMPUTATIONS: For $g=0,1,\dots,q-1$ recursively compute the values
@@
Alg. 7.1 outputs $\sigma$-soft inclusion correctly by virtue of Lemma 3.1.
Next we will prove that it also correctly claims exclusion.
@@
If Alg. 7.1 outputs exclusion, then $1+\frac{d}{v_{g}}>\sigma^{h}$ for $g=0,1,\dots,q-1$. Hence $\frac{1}{v_{g}}>2\sqrt{q}$ because $\sigma^{h}>1+2d\sqrt{q}$ according to initialization of Alg. 7.1. Therefore, the assumptions of Prop. 7.1 hold for $p_{h}(x)$ replacing $p(x)$, and hence the disc $D(0,1)$ contains no zeros of $p_{h}(x)$, but then it contains no zeros of $p(x)$ as well (see Observation 6.2 (iii)), and so exclusion is correct.
@@
Alg. 7.1 is a $\sigma$-soft e/i test, certifying inclusion at the cost of evaluation of polynomial $p(x)$ at $2hj$ points $x$ or exclusion at the cost of evaluation at $2hq$ points $x$ for three integers $h$, $j$, and $q$ such that $1\leq j\leq q$, $q>m$, and $h:=\lceil\log_{\sigma}(1+2d\sqrt{q})\rceil$, $h=O(\log(d))$ for $\sigma-1$ exceeding a positive constant.
@@ 7.3 A faster deterministic e/i test
Next we supply relevant estimates.
@@
Prop. 6.2, for $p_{h}$ replacing $p$, and the assumed bound $\frac{d-m}{\theta^{h}-1}\leq\frac{1}{6\sqrt{q}}$ together imply
@@
Alg. 7.2 claims inclusion if $v_{g}\geq\frac{1}{3\sqrt{q}}$ for some $g$, but then
@@
Now Observation 6.2 (iii) implies that Alg. 7.2 claims inclusion correctly.
Next we prove that it also claims exclusion correctly.
@@
Combine these bounds to obtain $\max_{g=0}^{q-1}\Big{|}\frac{f^{\prime}(\zeta_{q}^{g})}{f(\zeta_{q}^{g})}\Big{|}<\frac{1}{2\sqrt{q}}$. To complete the proof, apply Prop. 7.1 for $f(x)$ replacing $p(x)$ and for $m$ replacing $d$. ∎
Readily extend Prop. 7.2 to Alg. 7.2 by decreasing its bounds on $q$ and $h$ provided that #$\#(D(0,\theta))=m$ for $\theta\geq 1+6(d-m)\sqrt{q}$.
@@
Alg. 7.2 is a $\sigma$-soft e/i test: it certifies either inclusion or exclusion at the cost of evaluation of $p(x)$ at $2hj$ or $2hq$ points $x$, respectively, for three integers $h$, $j$, and $q$ such that $1\leq j\leq q$, $q>m:=\#(D(0,\theta)),~{}\sigma^{h}>1+6m\sqrt{q}$, $\theta\geq 1+6(d-m)\sqrt{q}$, and $h:=\lceil\log_{\sigma}(1+6m\sqrt{q})\rceil$, so that $h=O(\log(m))$ if $\sigma-1$ exceeds a positive constant, and in this case we only need $O(m\log(m)$ queries of the oracle to perform an e/i test.
@@ 7.4 A randomized e/i test
7.4.1. Introductory comments. Would Algs. 7.1 and 7.2 stay correct if we apply them at fewer evaluation points? They have consistently stayed correct for $q$ of order $\log^{2}(d)$ in extensive numerical tests of [45, 48] applied to $p(x)$ rather than to $p_{i}(x)$ for $i>0$. The tests reported exclusion where the values $|s_{h,q}|$ were reasonably small for $s_{h,q}:=\frac{1}{q}\sum_{g=0}^{q-1}\zeta^{(h+1)g}~{}\frac{p^{\prime}(\zeta^{g})}{p(\zeta^{g})},~{}{\rm for}~{}h=0,1,2$ (cf. Eqn. 12.2 in Part III), and reported inclusion otherwise. Clearly, $|s_{h,q}|\leq\delta$ for all $h$ if $|p(\zeta^{g})|\leq\delta/q$ for $g=0,1,\dots,q-1$, while inclusion is certified unless $|p(\zeta^{g})|$ is small for all $g$.
@@
Clearly, [45, 48] have never encountered worst case inputs in their extensive tests. Next, in line with this observation, we prove that our e/i tests, which evaluate $p(x)$ at $O(\log(d)$ points, are correct whp under the Random Root Model.
@@
Hence $|z-\bar{z}|=O(1/\beta^{g})$ for a fixed $\beta>1$. Combine this bound with (7.6) and (7.7), choosing sufficiently large $g$ of order $\log(d)$ unless $\sigma|\bar{z}|<\beta$.
@@
For a fixed $\sigma>1$, e.g., for $\sigma=1.2$, choose a constant $\beta>1$ and a sufficiently large $g=O(\log(d))$ such that $\gamma<\pi$ in (7.4). Then under the Random Root Model the probability $P$ of the output error for Alg. 3.1(rrm) is less than a constant $\nabla<1$. We can decrease the error probability below $\nabla^{v}$ by applying the algorithm $v$ times, in particular below $d^{a}$ for any fixed positive constant $a$ and for $v=O(\log(d))$. We arrive at the following
@@
Under Random Root Model one only needs to query an evaluation oracle to perform an e/i test correct;y whp.
@@
It is tempting to combine root-lifting or root-squaring with the algorithms of Part III to obtain an alternative randomized e/i test supporting Cor. 7.1, but for that we would need to raise disc isolation from order of $1+1/m$ to a constant; this would require too costly lifting of order $m$.